how many milliliters of water must be added to 50 ml of a 24% acid solution to make a solution that is 12% acid solution
Answer by checkley79(3341)

.24*50=.12(50+X)
12=6+.12X
.12X=12-6
.12X=6
X=6/.12
X=50 ML OF WATER IS NEEDED.
PROOF:
.24*50=.12(50+50)
12=.12*100
12=12